# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 5.6. Subtask alignment
To improve outcome quality at each step, align subtasks to worker capabilities (Figure 10).
Title. Subtask alignment, To improve outcome quality at each step, align subtasks to worker capabilities. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increased effort to design). Section 2: tactics used. Subtasks (partition, focus), Architectures (dynamic, communicative).
Motivation. One critical strategy is ensuring that subtasks match the workers’ capabilities appropriately. Simplifying tasks can increase outcome quality (Chilton et al., 2013; Wu et al., 2022b; Lasecki et al., 2013; Pietrowicz et al., 2013).
Implementation. Compared to entire tasks, smaller subtasks are often easier to complete, validate, and correct (Ambati et al., 2012; Schick et al., 2022; Wu et al., 2022b; Yang et al., 2022; Kim et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022). Designers can also adapt the subtask content to fit the worker’s capabilities (Chang et al., 2017; Mohanty et al., 2019; Pietrowicz et al., 2013; Arora et al., 2022). For example, researchers found that crowdworkers were better at generating predictive features than at estimating if a feature is predictive; so they adapted the workflow accordingly (Cheng and Bernstein, 2015).
When the appropriate subtasks are unknown, dynamic architectures can use partition or focus subtasks to use workers’ input to break down tasks into an appropriate size at runtime (Kittur et al., 2011; LaToza et al., 2014; Kim et al., 2017; Huang et al., 2023c). Communicative architectures can create appropriate subtasks for workers when they have a running “todo” list from which workers can choose their preferred task of many (Zhang et al., 2012; Lasecki et al., 2015; Mahyar et al., 2018).
Consequences. Much remains unknown about creating appropriate subtasks for LLMs. The capabilities of LLMs and appropriate subtask content are far from fully understood. Furthermore, it is unknown the degree to which LLMs can judge the appropriateness of subtasks for other LLMs to complete, bringing into question the effectiveness of some partition subtask and communicative architecture approaches taken when the workers are crowdworkers. Therefore, testing out these different options to find appropriate subtasks is resource intensive in the design process.
Examples. Examples that pay particular attention to subtask alignment include: Revolt (Chang et al., 2017), CrowdBand (Pietrowicz et al., 2013), and Ask Me Anything (Arora et al., 2022).
Related strategies. Subtasks must align to the capabilities of recruited workers and requirements must be communicated with sufficient context. Subtask alignment is especially challenging when subtasks are created during runtime eith adaptable architectures.
@@ 5.7. Context
To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of worker instructions (Figure 11).
Title. Context, To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of worker instructions. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Critical for quality outputs, Excessive context can confuse), Constraints (Increase effort to design), Accuracy (Affected by small changes). Section 2: tactics used. Subtasks (all), Architectures (communicative).
@@
Implementation. Communicating context to workers can help with improving the quality of the output and creating more coherence among disparate subtasks (Luther et al., 2015; Wu et al., 2022a; Teevan et al., 2016; Chilton et al., 2019; Hahn et al., 2016; Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010). There are many forms of contextual communication, from input text (Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010), to workflow goal tracking (Mirowski et al., 2023; Yang et al., 2022; Kim et al., 2017), to explicit communication between users and workers (Salehi et al., 2017).
@@
Instruction wording is a critical part of all subtask designs to support high outcome quality (Parameswaran et al., 2023). A key difficulty is communicating the user’s intent through language. One study reports that crowdworkers tend to need less explicitly defined instruction than LLMs, but can still struggle with vague or ambiguous instructions (Wu et al., 2023c). Both crowdworkers and LLMs benefit from examples, namely screening and tutorial tasks for crowdworkers (Teevan et al., 2016; Lasecki et al., 2015; Chen et al., 2019; Lasecki et al., 2013; Pietrowicz et al., 2013; Drapeau et al., 2016; Bernstein et al., 2010) and few-shot examples for LLMs (Nair et al., 2023a; Gero et al., 2023). Interfaces for crowdworkers can structure their output, while this structure needs to be communicated to LLMs in natural language (Wu et al., 2023c; Creswell and Shanahan, 2022; Kim et al., 2014).
@@
Consequences. Although context is generally important, overcomplicating instructions can lead to confusion (Chen et al., 2019; Lasecki et al., 2015; Wu et al., 2023c; Madaan et al., 2023b), so understanding the right amount of context to give is important (Kittur et al., 2011). One in-depth crowdsourcing study on providing context through interactions with the user found that providing context was helpful early on in the workflow if content is poor, but became detrimental if communicated later or with high-quality content (Salehi et al., 2017). Minimizing context seen can also help with preserving privacy (Bragg et al., 2021; Kaur et al., 2017). For LLMs, context may be included as part of the prompt, although LLMs have the additional barrier of an input token limit (Park et al., 2022; Mirowski et al., 2023). Therefore, workflow designs should aim for a minimum amount of sufficient context, to provide coherence while avoiding confusion.
@@
For both crowdworkers and LLMs, small changes in instructions or prompts can lead to dramatically different outcomes, impacting the accuracy of the final result (Wu et al., 2022b; Kittur et al., 2012; McDonnell et al., 2016; Wu et al., 2022a; Mirowski et al., 2023; Arora et al., 2022; Huang et al., 2023c; Wu et al., 2023c; Parameswaran et al., 2023; Gero et al., 2023). Additionally, different LLMs may respond differently to the same input prompt (Parameswaran et al., 2023). Therefore, figuring out the clearest instructions requires nontrivial effort and testing.
@@
Examples. Examples that pay particular attention to context include Turkomatic (Kulkarni et al., 2012), collaborative writing with microtasks (Teevan et al., 2016), and a study on communicating context (Salehi et al., 2017).
@@
Title. Quality thresholds, to more precisely balance the quality-cost tradeoff, measure quality repeatedly and stop the workflow once a quality threshold is reached. Section 1: effects on objectives. Several objectives listed (with their effects). Quality and Constraints (Improves control of tradeoff). Section 2: tactics used. Workers (LLMs, crowdworkers, user), Subtasks (evaluate), Architectures (redundant, dynamic).
@@ 5.8. Quality thresholds
Implementation. Some workflows give the worker control over the number of redundant generations using evaluate subtasks (Yao et al., 2023; Huffaker et al., 2020; Drapeau et al., 2016; Parameswaran et al., 2023). Others allow for tuning redundant architectures giving flexible goal precision, recall, and agreement thresholds (Huffaker et al., 2020; Yao et al., 2023), use decision-theoretic choices of the number of responses to obtain in redundant workflows (Bragg et al., 2013; Alshaibani and Quinn, 2021; Dai et al., 2010, 2013), or dynamically choose between workflows by their maximum expected value of utility (Goto et al., 2016).
@@
Examples. Examples of this strategy include POMDP-based control (Dai et al., 2013), Pterodactyl (Alshaibani and Quinn, 2021), and a workflow for detecting emotionally manipulative language (Huffaker et al., 2020).
@@
A line graph showing target number of categories on the X axis from 2-20, and The percentage error on the y axis from -1 to 1, in which 0% error means exactly hitting the target. Two lines are shown. One line, for Cascade-LLM, stays much closer to the 0% error line and has a narrower 95% confidence interval than the second line, zero-shot-target. At 11 or more categories, zero-shot-target has a negative percentage error rate.
@@ 6. Case studies
We choose three popular crowdsourcing tasks that incorporate different degrees of the accuracy and creativity design space objectives. We implement Cascade (Chilton et al., 2013) to build taxonomies with manipulable structures, Soylent (Bernstein et al., 2010) to shorten text to a variety of target lengths, and Mechanical Novel (Kim et al., 2017) to control the elements of a story. Diagrams of the workflows, as well as our adaptations to use LLMs, can be found in Figures 13, 16, and 18.

### v4

@@ 5.6. Subtask alignment
To improve outcome quality at each step, align subtasks to actor capabilities (Figure 11).
Title. Subtask alignment, To improve outcome quality at each step, align subtasks to actor capabilities. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increased effort to design). Section 2: tactics used. Subtasks (partition, focus), Architectures (dynamic, communicative).
Motivation. One critical strategy is ensuring that subtasks match the actors’ capabilities appropriately. Simplifying tasks can increase outcome quality [C(Lasecki et al., 2013; Chilton et al., 2013; Pietrowicz et al., 2013) L(Wu et al., 2021)].
Implementation. Compared to entire tasks, smaller subtasks are often easier to complete, validate, and correct [C(Ambati et al., 2012) L(Creswell et al., 2022; Wu et al., 2021; Creswell and Shanahan, 2022; Yang et al., 2022; Kim et al., 2023b; Schick et al., 2022)]. Designers can also adapt the subtask content to fit the actor’s capabilities [C(Pietrowicz et al., 2013; Mohanty et al., 2019; Chang et al., 2017) L(Arora et al., 2023)]. For example, researchers found that crowdworkers were better at generating predictive features than at estimating if a feature is predictive; so they adapted the workflow accordingly [C(Cheng and Bernstein, 2015)].
When the appropriate subtasks are unknown, dynamic architectures can use partition or focus subtasks to use actors’ input to break down tasks into an appropriate size at runtime [C(Kittur et al., 2011; LaToza et al., 2014; Kim et al., 2017) L(Huang et al., 2023c)]. Communicative architectures can create appropriate subtasks for actors when they have a running “todo” list from which actors can choose their preferred task of many [C(Mahyar et al., 2018; Zhang et al., 2012; Lasecki et al., 2015)].
Consequences. The capabilities of LLMs and appropriate subtask content are far from fully understood. Furthermore, it is unknown the degree to which LLMs can judge the appropriateness of subtasks for other LLMs to complete, bringing into question the effectiveness of some partition subtask and communicative architecture approaches taken when the actors are crowdworkers. Therefore, testing out these different options to find appropriate subtasks is resource intensive in the design process.
Examples. Examples that pay particular attention to subtask alignment include: Revolt [C(Chang et al., 2017)], CrowdBand [C(Pietrowicz et al., 2013)], and Ask Me Anything [L(Arora et al., 2023)].
Related strategies. Subtasks must align to the capabilities of selected actors and requirements must be communicated with sufficient context. Subtask alignment is especially challenging when subtasks are created during runtime with adaptable architectures.
@@ 5.7. Context
To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of actor instructions (Figure 12).
Title. Context, To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of actor instructions. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Critical for quality outputs, Excessive context can confuse), Constraints (Increase effort to design), Objectivity (Affected by small changes). Section 2: tactics used. Subtasks (all), Architectures (communicative).
@@
Implementation. Communicating context to workers can help with improving the quality of the output and creating more coherence among disparate subtasks [C(Luther et al., 2015; Chilton et al., 2019; Hahn et al., 2016; Teevan et al., 2016; Kulkarni et al., 2012; Bernstein et al., 2010; Ambati et al., 2012) L(Wu et al., 2022a)]. There are many forms of contextual communication, from input text [C(Kulkarni et al., 2012; Ambati et al., 2012; Bernstein et al., 2010)], to workflow goal tracking [C(Kim et al., 2017) L(Yang et al., 2022; Mirowski et al., 2023)], to explicit communication between users and workers [C(Salehi et al., 2017)].
@@
Instruction wording is a critical part of all subtask designs to support high outcome quality [L(Parameswaran et al., 2023)]. A key difficulty is communicating the user’s intent through language. One study reports that crowdworkers tend to need less explicitly defined instruction than LLMs but can still struggle with vague or ambiguous instructions [L(Wu et al., 2023b)]. Both crowdworkers and LLMs benefit from examples, namely screening and tutorial tasks for crowdworkers [C(Pietrowicz et al., 2013; Lasecki et al., 2015, 2013; Teevan et al., 2016; Drapeau et al., 2016; Bernstein et al., 2010; Chen et al., 2019)] and few-shot examples for LLMs [L(Gero et al., 2023; Nair et al., 2023a)]. Interfaces for crowdworkers can structure their output, while this structure needs to be communicated to LLMs in natural language [C(Kim et al., 2014) L(Wu et al., 2023b; Creswell and Shanahan, 2022)].
@@
Consequences. Although context is generally important, overcomplicating instructions can lead to confusion [C(Lasecki et al., 2015; Chen et al., 2019) L(Madaan et al., 2024; Wu et al., 2023b)], so giving the right amount of context is important [C(Kittur et al., 2011)]. One in-depth crowdsourcing study found that providing context from the user was helpful early on in the workflow if content is poor but becomes detrimental if communicated later or with high-quality content [C(Salehi et al., 2017)]. Minimizing context seen can also help with preserving privacy [C(Bragg et al., 2021; Kaur et al., 2017)]. For LLMs, context may be included as part of the prompt, although LLMs have the additional barrier of an input token limit [L(Mirowski et al., 2023; Park et al., 2022)]. Therefore, workflow designs should aim for a minimum amount of sufficient context, to provide coherence while avoiding confusion.
@@
For both crowdworkers and LLMs, small changes in instructions or prompts can lead to dramatically different outcomes, impacting the accuracy of the final result [C(McDonnell et al., 2016; Kittur et al., 2012) L(Wu et al., 2021; Parameswaran et al., 2023; Wu et al., 2023b; Gero et al., 2023; Wu et al., 2022a; Mirowski et al., 2023; Arora et al., 2023; Huang et al., 2023c)]. Additionally, different LLMs may respond differently to the same input prompt [L(Parameswaran et al., 2023)]. Therefore, figuring out the clearest instructions requires nontrivial effort and testing.
@@
Examples. Examples that pay particular attention to context include Turkomatic [C(Kulkarni et al., 2012)], collaborative writing with microtasks [C(Teevan et al., 2016)], and a study on communicating context [C(Salehi et al., 2017)].
@@
Title. Quality thresholds, to more precisely balance the quality-cost tradeoff, measure quality repeatedly and stop the workflow once a quality threshold is reached. Section 1: effects on objectives. Several objectives listed (with their effects). Quality and Constraints (Improves control of tradeoff). Section 2: tactics used. Actors (LLMs, crowdworkers, user), Subtasks (evaluate), Architectures (redundant, dynamic).
@@ 5.8. Quality thresholds
Implementation. Some workflows give the actor control over the number of redundant generations using evaluate subtasks [C(Huffaker et al., 2020; Drapeau et al., 2016) L(Parameswaran et al., 2023; Yao et al., 2024)]. Others allow for tuning redundant architectures giving flexible goal precision, recall, and agreement thresholds [C(Huffaker et al., 2020) L(Yao et al., 2024)], use decision-theoretic choices of the number of responses to obtain in redundant workflows [C(Bragg et al., 2013; Dai et al., 2010; Alshaibani and Quinn, 2021; Dai et al., 2013)], or dynamically choose between workflows by their maximum expected value of utility [C(Goto et al., 2016)].
@@
Examples. Examples of this strategy include POMDP-based control [C(Dai et al., 2013)], Pterodactyl [C(Alshaibani and Quinn, 2021)], and a workflow for detecting emotionally manipulative language [C(Huffaker et al., 2020)].
@@
[…]
6. Using the design space and strategies
In this section, we show the generative value of the design space through two vignettes. In contrast to the paper’s case studies (Section 7), which originate with an existing crowdsourcing workflow, here we use the design space to generate the designs of novel workflows from scratch.
After defining our example task of Retrieval-Augmented Generation (Section 6.1), we present two vignettes that implement this task in different contexts (Sections 6.2 and 6.3, Figure 14). These vignettes exemplify how our design space informs which tactic to employ in response to shifting task contexts. Section 6.4 discusses the function of the design space within the broader design process. While we have neither implemented nor evaluated these chains — and so make no claims that they perform better than existing approaches reported in the literature — they are (to our knowledge) novel and offer the potential for high performance.
6.1. Task: Retrieval Augmented Generation
We consider the task of Retrieval Augmented Generation (RAG), in which an LLM generates responses with the help of a larger corpus from which it can retrieve information. RAG systems are usually implemented with a multi-step chaining process due to context size (Shao et al., 2024). Specifically, we consider a RAG Question-Answering task in which the system must answer questions using information from the corpus.
Conceptually, we can break the RAG QA task into retrieval and generation steps. The retrieval step inputs the question and the corpus and outputs relevant information from the corpus. The generation step inputs the question and said relevant information, then outputs an answer. The recommendations provided by the LLM chaining software LangChain reflect the standard approach for RAG (Lan, [n. d.]). They recommend chunking and indexing documents in the corpus, using the query to retrieve documents by semantic similarity, then placing the retrieved chunks into an LLM context and generating an answer (Figure 14).
6.2. Vignette 1: Customer service
For this vignette, we consider implementing RAG as a part of a customer service system, intended to answer questions about company policies and terms of service.
Prioritized objective: objectivity
For customer queries, we will make the assumption that there is just one answer as all information in the corpus should be self-consistent. The diverse responses and validation strategies are especially useful for improving accuracy on objective tasks. Evaluate and improve subtasks support validation. Additionally, requiring a high threshold of agreement across redundant outputs, with a plan for reconciliation if outputs do not reach that standard, uses the diverse responses strategy to validate by measuring consistency.
Prioritized objective: task specific
This application of RAG is specific to one corpus about the company policies. The most helpful strategy for this objective is precise validation, since the specificity of the task allows for pre-definition of qualities to be validated. The tactical instantiation of this strategy is highly specific evaluate subtasks.
Prioritized objective: resource constraints
Latency and user effort lead to customer frustration so are important to minimize in a customer service system. Expense and effort to design are also desirable to minimize, but less critical in this task context. To reduce the strain on the latency of the diverse responses and validation strategies, the architecture should be a parallel redundant one, rather than iterative or sequential. Spending effort selecting actors, designing context, and aligning subtasks reduces latency, expense, and user effort with less impact on quality. Although the adaptable architectures and quality thresholds strategies can reduce latency, expense, and effort to design, they can induce quality issues best mitigated through user guidance. As reducing user effort is prioritized, this strategy should be deprioritized.
Chain design
An objectives-driven chain for this customer service task pieces together the various suggested tactics into a cohesive chain such as the following. After indexing and retrieval, the selected chunks are fed into the generation portion of the chain. This portion has multiple redundant generate subtasks to answer the query based on the retrieved information. They are then passed to redundant evaluate subtasks, in which the highest scoring generation provides the output. The number and subject of these evaluations should be determined through testing. Finally, if all outputs do not meet an agreement threshold, alternative outcomes are provided such as requesting a rewording of the query or connecting to a representative. See Figure 14 for a visual representation.
6.3. Vignette 2: Fitness corpus
For the second vignette, we consider a RAG system informed by a corpus of scientific papers to answer a wide variety of questions around fitness topics, such as exercise, nutrition, and sleep.
Prioritized objective: subjectivity
Objectivity in providing correct answers continues to be critical, but this application must incorporate uncertainty and open-endedness more than the first vignette, as the corpus of literature can support differing conclusions. The strategies of diverse responses, validation, and user guidance are especially helpful in addressing this uncertainty. To implement these strategies, the user can provide input at runtime to validate what output is desirable when there is uncertainty in the answer. Measuring uncertainty requires sourcing diverse responses, either through measuring variation within a redundant architecture or measuring ease of persuasion in the debate communicative architecture.
Prioritized objective: task general
The RAG system must generalize to a wider variety of queries and topics than the first vignette. Adaptable architectures provide the flexibility to support such variety. As the variety of queries is not pre-defined, branching architectures are not a useful tactical instantiation of this strategy. Instead, dynamic and communicative architectures can adapt at runtime to an unknown set of queries and topics. Finally, incorporating user guidance supports flexibility in result interpretation and reduces the need to specify a pre-defined set of topics or queries.
Prioritized objective: quality
Although quality is always desirable, in high-stakes domains like the medical field prioritizing this objective is critical even at higher expense, latency, or effort. For the purposes of this vignette, we assume few resource constraints. Of all the objectives, prioritizing an increase in quality has the widest range of options, as nearly all strategies and their tactical implementations improve quality. The exception is that adaptable architectures can induce errors and reduce transparency and consistency. In this health querying system, the user needs appropriate context to ensure the qualities of transparency and faithfulness. The chain’s tactics can attribute outputs to verbatim sections of the corpus to improve the user’s context.
Chain design
By combining these prioritized tactics, an objectives-driven chain for this case could look like the following. The initial retrieval step allows for multiple chunks of the corpus to pass through. To accommodate for the potential disagreement from a variety of sources, the next step iteratively merges the corpus chunks into several distinct lines of argument, connected with references to verbatim parts of the corpus. At each step of this iterative redundant architecture, a pairwise comparison evaluate subtask determines if two chunks support similar takeaways. If they do, a subsequent subtask merges the two into a new chunk, collecting the indexes referencing specific parts of the corpus. This iterative loop runs for a dynamically determined amount of time until the arguments have converged. These arguments and their references then pass into the generation portion of the pipeline. Each argument informs a generation subtask that answers the query given that set of retrieved information. Then, a communicative debate commences in which each debater begins with this generated answer and its justification. Using the ease with which each debater is convinced it is incorrect, the chain outputs the different arguments, their uncertainty, the debate history, and references to the initial corpus for the user to evaluate. See Figure 14 for a visual representation of this chain.
6.4. Vignette discussion
Through these two vignettes, we outline how the design space specializes chains to the task context by explicitly defining objectives and surfacing the cascading implications for strategies that then recommend beneficial tactic choices. In this section, we discuss how the process described in these vignettes fits into the broader design process.
A line graph showing the target number of categories on the X-axis from 2-20, and The percentage error on the y-axis from -1 to 1, in which 0% error means exactly hitting the target. Two lines are shown. One line, for Cascade-LLM, stays much closer to the 0% error line and has a narrower 95% confidence interval than the second line, zero-shot-target. At 11 or more categories, zero-shot-target has a negative percentage error rate.
Tactic suggestions respond to task context.
Standard practice applies the same steps for any task context: retrieval and generation. With our design space, a designer can flexibly accommodate unique combinations of objectives for varied applications, as seen in the differing vignette designs.
Incorporating non-LLM tools can reduce costs.
The costs of using LLMs for all steps can be prohibitive and unnecessary. A designer could replace the semantic similarity retrieval step with LLMs by chunking the corpus and using focus subtasks to determine relevant parts of each chunk. Although the potential for LLMs for processing complex unstructured data is an active area of exploration (Shankar et al., 2024), using LLMs for this step would cost much more than standard retrieval with a large corpus. Our design space focuses on the LLM chaining literature, but there is immense opportunity and future work in incorporating tools beyond LLMs into these chains.
@@ 6.4. Vignette discussion
The design space filters and inspires, rather than prescribes.
Using the design space does not prescribe a definitively optimal chain, but it does serve the design process through filtering and generating ideas. Through the vignettes, we see that the design space prioritizes subsets of options within a task context. Given the enormous space of possible designs, this filtering is valuable guidance. Furthermore, using the design space allowed us to create unique chains. Due to the comprehensiveness of the design space over two literatures, it serves as inspiration during the brainstorming process for a variety of potentially unintuitive options validated through prior work.
The design space provides an informed starting point, rather than a finished design.
The design space serves as a launching pad for other design decisions. Further testing can compare performance within the subset of tactic choices recommended by the design space (e.g., which type of verification performs best). Many of these changes depend on model performance (e.g., are diverse generations necessary, or is fluency high enough to have confidence in a single generation). Additionally, the design process will need to test specific tactical design decisions (e.g., what model to choose, prompt engineering, how many iterations, and for which specific qualities to evaluate). Although these chain designs may update after testing on specific corpora and models, they provide a starting point motivated by findings from prior work. Future work is needed to improve the granularity of design choice recommendations and develop evaluation mechanisms for comparing among design choices (Section 9.4).
@@ 7. Case studies
We choose three popular crowdsourcing tasks that incorporate different degrees of the objectivity and subjectivity design space objectives. We implement Cascade (Chilton et al., 2013) to build taxonomies with manipulable structures, Soylent (Bernstein et al., 2010) to shorten text to a variety of target lengths, and Mechanical Novel (Kim et al., 2017) to control the elements of a story. Diagrams of the workflows, as well as our adaptations to use LLMs, can be found in Figures 16, 18, and 20.
@@
Four taxonomies. The first with default parameters (merge threshold 0.75, minimum category size 2), contains a hierarchy of 4 layers and 23 categories. The second that is more likely to merge (merge threshold 0.5, min category size 2), contains 4 layers and 18 categories. The third is a larger minimum category size (merge threshold 0.75, min category size 5) contains 3 layers and 13 categories). The fourth is both more likely to merge and has a larger min category size (merge threshold 0.5, min category size 5), contains 3 layers and 9 categories).
7.1. Case study 1: Taxonomy creation
For our first case study, we explore taxonomy creation, the task of organizing a set of items into a category tree to aid sensemaking. This task’s outcome qualities are primarily objective, with some degree of subjectivity: categories need to be correct, but most items can fit into multiple categorizations of different levels of precision. We focus on the ability to control the granularity of these categorizations by controlling the taxonomy shape (Figure 17).
7.1.1. Workflow design
To build Cascade-LLM, our version of Cascade powered by LLMs, we made multiple changes to the original workflow. We highlight some of the most salient modifications. To achieve diverse categorizations, we improved response diversity by giving different prompts to parallel generations and ensuring that every item is input into a category generation step. At first, the categories were too broad. Specifying category qualities via targeted prompts proved unsuccessful, so instead we added a programmatic validation step to delete categories that included “and,” “or,” commas, or slashes. As this hard-coded validation step worked, it was not necessary to include another LLM call. The original workflow gives actors sets of eight items for which to generate eight categories. However, the LLM would try categorizing multiple items at once, consistent with prior findings that LLMs struggle to disambiguate multiple subtasks in a prompt (Wu et al., 2023b; Madaan et al., 2024). To increase diversity, we aligned subtask design to fit model capabilities with one item categorization. Finally, the original workflow also requires parameters for the minimum category size and a merge threshold. To support user guidance through direct manipulation, we dynamically apply these parameters at the last taxonomization stage. To investigate different taxonomies with varied levels of precision, a user can manipulate the parameters of 1) the minimum category size and 2) the similarity threshold determining if two categories will merge (Figure 17).
7.1.2. Evaluation setup